FT MICROWAVE STRUCTURE AND ANALYSIS OF FOUR ISOTOPOMERS OF THE ARGON-THIETANE VAN DER WAALS COMPLEX

JESSICA SCHLIER, STEWART E. NOVICK AND \underline{W. C. PRINGLE}, Department of Chemistry, Wesleyan University, Middletown, CT 06459; D. McCAMANT, Department of Chemistry, University of California , Berkeley, CA 94720.

The Fourier Transform Pulsed Jet Microwave spectra of the argon-thietane (C_{4}H_{6}S) van der Waals complex has been assigned for four isotopomers including ^{32}S and ^{34}S and each of the ^{13}C ^{32}S isotopomers. We previously reported analysis of the parent spectrum^{a}. The argon freezes out the ring inversion and converts the inversion dipole moment to a permanent moment. Using the isotopomers yields the unambiguous position of the argon that could not be determined from the parent complex alone. The argon is 3.780 Å above the plane containing the C-S-C bond and is 0.576 Å across the ring from the plane perpendicular to this angle containing the bisector of C-S-C angle. The CH_{2} across the ring from the sulfur is below the C-S-C plane. This latter position could only be determined using isotopomers. Large amplitude cross ring argon motion causes the two carbons bonded to the sulfur to be equivalent.

aW. C. PRINGLE, et. al., paper TD 12, presented at the 51st Ohio State University International Symposium on Molecular Spectroscopy, Columbus, OH, June 1996.